OpenStudy (anonymous):

Peter deposited $8,975 into a savings account 21 years ago. The account has an interest rate of 3.8% and the balance is currently $19,909.20. How often does the interest compound?

OpenStudy (anonymous):

daily monthly quarterly annually (answer choices)

OpenStudy (jamesj):

The easiest way to do this is to calculate for each option until you find an answer that matches the final balance. For example, if it compounded annually, then the amount would be \[ 8975(1+.038)^21 \] Is that equal to 19,909.2?

OpenStudy (jamesj):

**correction: \( 8975(1+.038)^{21} \)

OpenStudy (anonymous):

so would it be daily monthly quarterly or annually?

OpenStudy (jamesj):

Keep calculating and see.

OpenStudy (anonymous):

i don't get it..

OpenStudy (jamesj):

If it money were compounded quarterly, then what is the expression for the value of $8,975 at the end of 21 years?

OpenStudy (anonymous):

it'd be yearly?

OpenStudy (jamesj):

I wrote down above the expression for annual or yearly compounding: \[ 8975(1+0.038)^{21} \] That is equal to $19,641.70, so that's not the answer. Now what is the expression if the amount is compounded quarterly?

OpenStudy (anonymous):

okay if its quaterly where do i put 24 at? like what do i multiply that by?

OpenStudy (jamesj):

What's the value of $1 at 3.8% compounded quarterly at the end of one year? It is \[ $1 . (1 + \frac{0.038}{4})^4 \] Make sense?

OpenStudy (jamesj):

If so, then the value of $8975 at 3.8% compounded quarterly at the end of 21 years is \[ $8975(1 + \frac{0.038}{4})^{21 \times 4 } \]

OpenStudy (anonymous):

ohhh okay lol